Please help me simplify X^1/2 divided by X^1/4 = ?
Answer by rothauserc(4718)

X^1/2 divided by X^1/4 = X^1/2 * X^-1/4
note that 1/X^1/4 = X^-1/4
= X^1/4
note that X^1/2 = X^2/4